An electric of 2 HP has an efficiency of 60%. If it works to lif water from a well of depth 20m for 10 min, find how much water will be left on the ground. (Take g=10m/s)
 
*Please explain what an HP is  and how to solve this problem. *

Grade:11

1 Answers

Vikas TU
7 years ago
HP stands  for Horse Power and
1 HP =746 watts.
since the efficiency is 60%
Power used to lift up the water => 0.6 * 2*746 = > 895.2 Watt.
P =W/t
   P = mgh/t 
   895.2  =  m(10)(10)/(10*60)
m = 5371.2 g or 5.3712 kg of water.